# Federated Sketching LoRA: On-Device Collaborative Fine-Tuning of Large Language Models

### v2

@@ 5.2 Ablation Study
Impact of sketching: In Figures 2 and LABEL:fig:rank_varying, we compare the performance of FSLoRA with and without sketching on fine-tuning the RoBERTa model and the LLaMA-3.2-3B model, respectively. Notably, FSLoRA without sketching is equivalent to the vanilla Federated LoRA. For FSLoRA with sketching, we apply a uniform sketching ratio of $k_{i}/r=0.5$ across all distributed devices. The upload budget for each device is set to $100$ and $200$ times the size of the full global LoRA modules at the corresponding rank for the RoBERTa and the LLaMA-3.2-3B models, respectively. As shown in Figures 2 and LABEL:fig:rank_varying, both FSLoRA with and without sketching achieve higher accuracy when the rank $r$ increases due to the availability of more tunable parameters. In addition, FSLoRA consistently outperforms its non-sketched counterpart across all the ranks and datasets. The use of sketching increases the communication frequency for devices under the same communication budget, thereby facilitating the optimization process and enhancing fine-tuning efficiency.
Figure 2: Comparison between FSLoRA with and without sketching (i.e., vanilla Federated LoRA), where the upload budget for devices is set to 100×100$\times$100 × the size of the global LoRA modules at each rank. FSLoRA with sketching obtains a better performance, validating its communication efficiency.
Impact of the global rank: In Figure LABEL:fig:global_rank, we investigate the impact of the rank of the global LoRA modules on FSLoRA’s performance. We vary the rank of the global LoRA modules while keeping the rank of submatrices updated by the devices to be consistent (i.e., $k_{i}=8$). This ensures that the communication and computational resources on the client side remain unchanged. As illustrated in Figure LABEL:fig:global_rank, FSLoRA maintains stable convergence across all the configurations. Moreover, FSLoRA demonstrates improved performance as the global rank increases. This observation confirms that the proposed sketching mechanism enables resource-constrained systems to reap the benefits of a higher global rank, striking an effective balance between efficiency and performance.
@@
Further experiments: Additional results, including detailed comparisons on each task in the commonsense reasoning benchmark corresponding to Figures LABEL:fig:rank_varying and LABEL:fig:global_rank, and the integration of compression and sketching, are provided in Appendix E. Results with more devices and broader heterogeneity, and a larger model can be found in Appendix C and Appendix D, respectively.
6 Conclusion and Limitation
We have proposed FSLoRA, a novel on-device collaborative LLM fine-tuning framework that introduces a sketching mechanism to enhance both performance and efficiency in resource-constrained systems. By maintaining large-rank LoRA modules on the server and allowing devices to selectively update submatrices based on the sketching ratios, FSLoRA effectively adapts to heterogeneous communication and computational constraints. We provide a rigorous convergence analysis of FSLoRA that characterizes how the sketching ratios affect the convergence rate. Finally, we confirmed the effectiveness of FSLoRA through extensive experiments across multiple datasets and models. A limitation of this work is the absence of a real-world deployment study that evaluates FSLoRA and baseline methods on the actual server and devices due to hardware constraints.
@@ Appendix A Comparison of Computation, Memory, and Communication
Computation and memory: Let $P$ and $q$ denote the memory cost of the full model and the global LoRA module (rank $r$), respectively. The computational cost is expressed with the big O notation. Forward and backward computations, as well as activation memory, are omitted as they are identical across all the considered methods. The results are summarized in Tables A.1 and A.2, where $m$ and $n$ denote the shape of the base model, $k_{i}$ denotes the LoRA rank for device $i$, $H$ denotes the number of iterations per round, and $N$ is the number of devices. Additionally, the results for the vanilla Federated LoRA, denoted as FedLoRA, are reported under the case of homogeneous LoRA ranks, i,e., $k_{i}=r$.
@@
FedStackLoRA $P+\max\left\{\sum_{i=1}^{N}\frac{k_{i}}{r}q,P\right\}$ $\mathcal{O}\left(Hk_{i}(m+n))+(\sum_{i=1}^{N}k_{i})mn+mn\right)$
@@
FedStackLoRA $\sum_{i=1}^{N}\frac{k_{i}}{r}q$ $\mathcal{O}\left((\sum_{i=1}^{N}k_{i})(m+n)\right)$
@@
As shown in Tables A.1 and A.2, FSLoRA matches HetLoRA in both computation and memory cost. FedStackLoRA introduces additional client-side overhead due to merging LoRA modules. FlexLoRA incurs extra server-side costs from conducting SVD on the full model. In summary, FSLoRA guarantees convergence with minimum overhead.
Communication: We detailed the communication load for baselines and our methods in Table A.3, where $q$ denotes the communication cost of a global LoRA module with rank $r$, $k_{i}$ denotes the local LoRA rank for device $i$, $m$ and $n$ denote the shape of the base model, and $N$ denotes the number of devices.
@@
FedLoRA HeteroLoRA FlexLoRA FedStackLoRA FSLoRA
@@
For the uplink, all four heterogeneous LoRA algorithms incur the same communication overhead for transmitting updated local LoRA modules, which is lower than that of FedLoRA. For the downlink, FedStackLoRA requires broadcasting the stacked LoRA modules, while HeteroLoRA and FlexLoRA broadcast the updated global LoRA modules. FSLoRA, on the other hand, broadcasts both the global LoRA modules and additional sketching matrices. The extra communication introduced by the sketching matrices is negligible compared to that of the global LoRA modules, as it consists only of binary sketching indices (i.e., the diagonal elements of the sketching matrix). For instance, in the case of the LLaMA-3.2-3B model under our experimental LoRA configuration, the global LoRA modules contain 66,060,288 parameters, equivalent to approximately 252 MB when using float32. With a global rank of $r=64$, the sketching indices require only 64 bits per device, covering all LoRA layers. Even with 100 devices, the total sketching overhead is merely 0.78 KB, which accounts for only 0.0003% of the global LoRA modules.
@@ Appendix B Difference between FSLoRA and FedBCGD
Both FSLoRA and federated block coordinate gradient descent (FedBCGD) [26] are motivated by device heterogeneity but are designed for fundamentally different deployment contexts. FedBCGD partitions the full model $\mathbf{x}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{N},\mathbf{x}_{s}]$, assigning each block $\mathbf{x}_{j}$ to a subset of devices with similar resource constraints, while the shared block $\mathbf{x}_{s}$ is optimized across all devices. While this block-partitioning strategy is effective for smaller models, it relies on explicit and static allocation, which can limit scalability and flexibility. As such, FedBCGD and similar block coordinate methods based on the full model are less suitable for on-device LLM fine-tuning.
FSLoRA, in contrast, builds on LoRA and introduces sparse diagonal sketching. Given a sketch matrix $\bf{S}$, the gradients of the loss $\ell({\bf W}_{0}+{\bf BSA},\xi)$ with respect to the LoRA matrices ${\bf B}$ and ${\bf A}$ are sparse: only selected columns of ${\bf B}$ and rows of ${\bf A}$ are updated in each round. By configuring the rank and sparsity of the sketch matrix $\bf{S}$, FSLoRA flexibly controls both the computational and communication load per device, enabling adaptation to heterogeneous device capabilities.
@@
Appendix C Evaluation under Broader Heterogeneity and Increased Number of Devices
C.1 Increasing Device Heterogeneity and the Number of Devices
@@
We extend our experiments on LLaMA-3.2-3B with the commonsense reasoning benchmark to $50$ devices. We adopt Dirichlet-based partitioning for dataset splits. Specifically, the commonsense reasoning benchmark includes $8$ tasks, and we partitioned them based on the Dirichlet distribution to construct task heterogeneity among $50$ devices. The Dirichlet concentration parameter is set to $\alpha=0.1$. We simulate device heterogeneity via different LoRA rank distributions (beyond the limited sketching ratio considered in Section 5). More capable devices are assigned higher ranks, reflecting varying compute capacities. We consider two different rank distributions: normal and heavy-tail distributions in the range $[4,64]$.
@@
Normal distribution: Ranks are sampled from a normal distribution with mean $\mu=\frac{a+b}{2}$ and standard deviation $\sigma=\frac{b-a}{6}$, where $a=4$ and $b=64$. This models a balanced distribution of device capabilities centered around the middle of the range.
@@
Heavy-tail distribution: We sample ranks using an inverse log-normal distribution. Specifically, we draw $x_{i}\sim\text{LogNormal}(\mu,\sigma)$ with $\mu=\log\left(\frac{a+b}{4}\right)$ and $\sigma=1.0$, then set $k_{i}=1/x_{i}$ and apply min-max normalization to scale values into the range $[a,b]$. This results in a heavy-tailed distribution where most devices receive low ranks, reflecting a scenario with many low-capability devices and a few high-capability ones.
@@
Table C.1: Accuracy comparison under different device heterogeneity settings. FSLoRA outperforms baseline methods across both normal and heavy-tail LoRA rank distributions.
@@
As shown in Table C.1, FSLoRA outperforms other methods under both heterogeneity settings. As we move from normal to heavy-tail, where more devices are low-resource, overall performance decreases for all methods. However, FSLoRA exhibits the smallest drop, demonstrating stronger robustness to extreme device heterogeneity.
In Figure 5, we compare the convergence behavior of FSLoRA and three baseline methods under the aforementioned two types of device heterogeneity. Under the normal distribution, FlexLoRA exhibits fast initial progress but falls behind FSLoRA in final accuracy, likely due to approximation errors introduced by truncated SVD. This issue is exacerbated in the heavy-tail distribution, where low-rank devices dominate and SVD truncation causes greater distortion, further degrading FlexLoRA’s performance. Similarly, HeteroLoRA’s reliance on zero-padding reduces optimization efficiency, preventing it from achieving higher accuracy. FedStackLoRA fails to show steady improvement as communication progresses. One potential reason is that frequent model merging and random reinitialization of LoRA modules in each round disrupt the convergence continuity. In contrast, FSLoRA demonstrates robust and stable convergence across both scenarios, achieving the highest overall accuracy.
@@
C.2 Further Increasing the Number of Devices
We further evaluated the performance of FSLoRA by increasing the number of devices to $100$. The results are presented in Table C.2. In this setting, local ranks follow a heavy-tailed distribution as described in the previous subsection, and all other experimental configurations remain unchanged. As shown in the table, FSLoRA maintains its advantage in terms of the average performance when scaling to more devices.
Table C.2: Accuracy comparison when the number of devices is $N=100$. FSLoRA maintains its advantage in terms of the average accuracy.

### v3

@@ 5.2 Ablation Study
Impact of sketching: In Figures 2 and LABEL:fig:rank_varying, we compare the performance of FSLoRA with and without sketching on fine-tuning the RoBERTa model and the LLaMA-3.2-3B model, respectively. Notably, FSLoRA without sketching is equivalent to the vanilla Federated LoRA. For FSLoRA with sketching, we apply a uniform sketching ratio of $k_{i}/r=0.5$ across all distributed clients. The upload budget for each client is set to $100$ and $80$ times the size of the full global LoRA modules at the corresponding rank for the RoBERTa and the LLaMA-3.2-3B models, respectively. As shown in Figures 2 and LABEL:fig:rank_varying, both FSLoRA with and without sketching achieve higher accuracy when the rank $r$ increases due to the availability of more tunable parameters. In addition, FSLoRA consistently outperforms its non-sketched counterpart across all the ranks and datasets. The use of sketching increases the communication frequency for clients under the same communication budget, thereby facilitating the optimization process and enhancing fine-tuning efficiency.
Figure 2: Comparison between FSLoRA with and without sketching (the latter equivalent to Federated LoRA) where the upload budget for clients is set to $100\times$ the size of the global LoRA modules at each rank. FSLoRA obtains a better performance, validating its communication efficiency.
(a) Fine-tuning the LLaMA-3.2-3B model on the commonsense reasoning benchmark. The results are averaged over eight tasks, illustrating FSLoRA’s ability to maintain strong performance while adapting to different rank and sketching configurations.
Impact of the global rank: In Figure LABEL:fig:global_rank, we investigate the impact of the rank of the global LoRA modules on FSLoRA’s performance. We vary the rank of the global LoRA modules while keeping the rank of submatrices updated by the clients to be consistent (i.e., $k_{i}=8$). This ensures that the communication and computational resources on the client side remain unchanged. As illustrated in Figure LABEL:fig:global_rank, FSLoRA maintains stable convergence across all the configurations. Moreover, FSLoRA demonstrates improved performance as the global rank increases. This observation confirms that the proposed sketching mechanism enables resource-constrained systems to reap the benefits of a higher global rank, striking an effective balance between efficiency and performance.
@@
Further experiments: Results with more clients under broader heterogeneity, as well as with a larger model, are reported in Appendix F and Appendix G, respectively. Appendix H.1 provides detailed per-task comparisons on the commonsense reasoning benchmark corresponding to Figures LABEL:fig:rank_varying and LABEL:fig:global_rank. The impact of varying the number of local updates $H$ is studied in Appendix H.2, while the extension to dynamic sketching ratios is presented in Appendix H.3. Finally, Appendix H.4 demonstrates the synergistic effect of integrating compression with sketching.
6 Conclusion
We have proposed FSLoRA, a novel collaborative LLM fine-tuning framework that introduces a sketching mechanism to enhance both performance and efficiency in resource-constrained systems. By maintaining large-rank LoRA modules on the server and allowing clients to selectively update submatrices based on the sketching ratios, FSLoRA effectively adapts to heterogeneous communication and computational constraints. We provide a rigorous convergence analysis of FSLoRA that characterizes how the sketching ratios affect the convergence rate. Finally, we confirmed the effectiveness of FSLoRA through extensive experiments across multiple datasets and models.
@@ Appendix A Comparison of Computation, Memory, and Communication
Computation and memory: Let $P$ and $q$ denote the memory cost of the full model and the global LoRA module (rank $r$), respectively. The computational cost is expressed with the big O notation. Forward and backward computations, as well as activation memory, are omitted as they are identical across all the considered methods. The results are summarized in Tables A.1 and A.2, where $m$ and $n$ denote the shape of the base model, $k_{i}$ denotes the LoRA rank for client $i$, $H$ denotes the number of iterations per round, and $N$ is the number of clients. Additionally, the results for the vanilla Federated LoRA, denoted as FedLoRA, are reported under the case of homogeneous LoRA ranks, i,e., $k_{i}=r$.
@@
FLoRA $P+\max\left\{\sum_{i=1}^{N}\frac{k_{i}}{r}q,P\right\}$ $\mathcal{O}\left(Hk_{i}(m+n))+(\sum_{i=1}^{N}k_{i})mn+mn\right)$
@@
FLoRA $\sum_{i=1}^{N}\frac{k_{i}}{r}q$ $\mathcal{O}\left((\sum_{i=1}^{N}k_{i})(m+n)\right)$
@@
As shown in Tables A.1 and A.2, FSLoRA matches HetLoRA in both computation and memory cost. FLoRA introduces additional client-side overhead due to merging LoRA modules. FlexLoRA incurs extra server-side costs from conducting SVD on the full model. In summary, FSLoRA guarantees convergence with minimum overhead.
Communication: We detailed the communication load for baselines and our methods in Table A.3, where $q$ denotes the communication cost of a global LoRA module with rank $r$, $k_{i}$ denotes the local LoRA rank for client $i$, $m$ and $n$ denote the shape of the base model, and $N$ denotes the number of clients.
@@
FedLoRA HeteroLoRA FlexLoRA FLoRA FSLoRA
@@
For the uplink, all four heterogeneous LoRA algorithms incur the same communication overhead for transmitting updated local LoRA modules, which is lower than that of FedLoRA. For the downlink, FLoRA requires broadcasting the stacked LoRA modules, while HeteroLoRA and FlexLoRA broadcast the updated global LoRA modules. FSLoRA, on the other hand, broadcasts both the global LoRA modules and additional sketching matrices. The extra communication introduced by the sketching matrices is negligible compared to that of the global LoRA modules, as it consists only of binary sketching indices (i.e., the diagonal elements of the sketching matrix). For instance, in the case of the LLaMA-3.2-3B model under our experimental LoRA configuration, the global LoRA modules contain 66,060,288 parameters, equivalent to approximately 252 MB when using float32. With a global rank of $r=64$, the sketching indices require only 64 bits per client, covering all LoRA layers. Even with 100 clients, the total sketching overhead is merely 0.78 KB, which accounts for only 0.0003% of the global LoRA modules.
@@ Appendix B Difference between FSLoRA and FedBCGD
Both FSLoRA and federated block coordinate gradient descent (FedBCGD) [28] are motivated by client heterogeneity but are designed for fundamentally different deployment contexts. FedBCGD partitions the full model $\mathbf{x}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{N},\mathbf{x}_{s}]$, assigning each block $\mathbf{x}_{j}$ to a subset of clients with similar resource constraints, while the shared block $\mathbf{x}_{s}$ is optimized across all clients. While this block-partitioning strategy is effective for smaller models, it relies on explicit and static allocation, which can limit scalability and flexibility. As such, FedBCGD and similar block coordinate methods based on the full model are less suitable for federated LLM fine-tuning.
FSLoRA, in contrast, builds on LoRA and introduces sparse diagonal sketching. Given a sketch matrix $\bf{S}$, the gradients of the loss $\ell({\bf W}_{0}+{\bf BSA},\xi)$ with respect to the LoRA matrices $\bf B$} and $\bf A$} are sparse: only selected columns of $\bf B$} and rows of $\bf A$} are updated in each round. By configuring the rank and sparsity of the sketch matrix $\bf{S}$, FSLoRA flexibly controls both the computational and communication load per client, enabling adaptation to heterogeneous client capabilities.
@@
Appendix C Justification for Random-$k$ Sketching
FSLoRA is built upon Random-$k$ diagonal sketching due to two key properties:
  • Submatrix selection. Given a sparse diagonal matrix $\mathbf{S}_{i}$, we have
    $\mathbf{BS}_{i}\mathbf{A}=\sum_{j\in\mathcal{I}_{i}}s_{j}\mathbf{b}_{j}\mathbf{a}_{j}^{\top}=[\mathbf{b}_{j}]_{j\in\mathcal{I}_{i}}\operatorname{diag}\{s_{j}\}_{j\in\mathcal{I}_{i}}[\mathbf{a}_{j}^{\top}]_{j\in\mathcal{I}_{i}},$
    where $\mathcal{I}_{i}$ corresponds to the index set of non-zero diagonal entries of $\mathbf{S}_{i}$ and $\mathbf{b}_{j}$ and $\mathbf{a}_{j}^{\top}$ denote the $j$-th column of module $\mathbf{B}$ and the $j$-th row of module $\mathbf{A}$, respectively. In other words, with Random-$k$ sketching, only a subset of $\mathbf{B}$’s columns and $\mathbf{A}$’s rows are activated for client $i$. The sparse diagonal structure effectively reduces local training cost for each client.
  • Unbiasedness for convergence. When $\mathbf{S}_{i}$ is a Random-$k$ diagonal matrix with $k_{i}$ nonzero diagonal entries $s_{j}=\frac{r}{k_{i}}$, we have
    $\mathbb{E}[\mathbf{BS}_{i}\mathbf{A}]=\mathbf{BA}.$
    This unbiasedness is critical for our convergence analysis.
Table C.1: Accuracy comparison of Random-$k$ sketching and importance-based sketching under the commonsense reasoning benchmark with the LLaMA-3.2-3B model. Random-$k$ sketching achieves better performance.
Importance metric ARC-c ARC-e BoolQ HSwag OBQA PIQA SIQA Wino Avg.
$\|a\|\|b\|$ 71.9 86.5 55.2 75.4 73.4 81.1 72.5 69.7 73.2
$\|a\|+\|b\|$ 72.1 86.4 64.5 76.8 70.8 82.2 71.3 69.3 74.2
[…]
In addition, we experimentally compare Random-$k$ sketching with importance-based sketching. For importance-based sketching, we sample (sketch) $k_{i}$ components from $\{\mathbf{b}_{j}\mathbf{a}_{j}^{\top}\}_{j=1}^{r}$ with probability set as the importance scores, e.g., $\|\mathbf{b}_{j}\|_{2}+\|\mathbf{a}_{j}\|_{2}$ or spectral norm $\|\mathbf{b}_{j}\|_{2}\cdot\|\mathbf{a}_{j}\|_{2}$. The results are shown in Table C.1. The results show that Random-$k$ diagonal sketching outperforms these choices for importance-based sketching. This may be because these importance measures are heuristic and do not reliably reflect actual contribution. Moreover, such sketching violates the unbiasedness property, complicating theoretical guarantees. While improved importance-based methods may enhance performance and remain a promising direction for future investigation, our current empirical and theoretical results favor random sketching.
Appendix D Compatibility of FSLoRA with Secure Aggregation
The aggregation of FSLoRA is compatible with secure aggregation. Taking the aggregation of module $\mathbf{B}$ as an example, we illustrate this below.
In FSLoRA, client updates are sparse matrices with non-zero values only in columns indexed by $\mathcal{I}_{i}\subset[r]$, size $|\mathcal{I}_{i}|=k_{i}$. With secure aggregation, each client apply additive masking:
$\tilde{\mathbf{B}}_{i}=\Delta\mathbf{B}_{i}+\mathbf{R}_{i},$
where mask $\mathbf{R}_{i}$ satisfies $\operatorname{supp}(\mathbf{R}_{i})\subseteq(u,v):u\in[m],v\in\mathcal{I}_{i}$ and $\sum_{i=1}^{N}\mathbf{R}_{i}=0$. That is, the mask has non-zero entries only in the client’s active columns, and all masks together sum to zero to preserve correctness. Such masks can be constructed following the classical protocol: for each pair of clients $(i,j)$, define a random matrix
$\mathbf{M}_{ij}=-\mathbf{M}_{ji}\in\mathbb{R}^{m\times r},\quad\operatorname{supp}(\mathbf{M}_{ij})\subseteq(u,v)\;|\;u\in[m],\,v\in\mathcal{I}_{i}\cap\mathcal{I}_{j},$
and then construct its total mask
$\mathbf{R}_{i}=\sum_{j>i}\mathbf{M}_{ij}-\sum_{j<i}\mathbf{M}_{ji}.$
During uploading, client $i$ sends the masked $k_{i}$ non-zero columns of $\tilde{\mathbf{B}}_{i}$, and then the server adds the corresponding padding and averages them as:
$\sum_{i=1}^{N}\tilde{\mathbf{B}}_{i}=\sum_{i=1}^{N}\left(\Delta\mathbf{B}_{i}+\mathbf{R}_{i}\right)=\sum_{i=1}^{N}\Delta\mathbf{B}_{i},$
which matches the aggregation of module $\mathbf{B}$ in (5). The aggregation of module $\mathbf{B}$ in FSLoRA is thus compatible with secure aggregation.
We can draw the same conclusion for module $\mathbf{A}$ under the same derivation.
Appendix E Empirical Validation of Assumptions
In the context of LLM fine-tuning, both the magnitude of stochastic gradients and gradients in LLM fine-tuning are in a mild range since:
  • Transformer-based LLMs use LayerNorm and scaled softmax attention, which stabilize activations and suppress gradient spikes.
  • Fine-tuning starts from a well-pretrained model already near a local minimum, leading to smaller gradients.
  • The fine-tuning dataset does not typically contain strong contradictory signals to what the model already knows, resulting in a relatively flat loss surface.
To further support this empirically, we report the statistics of the expected norm of the stochastic gradients $\mathbb{E}\|\nabla_{\mathbf{X}}\tilde{\ell}(\mathbf{X},\xi;\mathbf{S})\|$ over approximately $4500$ samples for 4 representative clients with different sketching ratios. The table below reports the minimum and maximum expected norm among 30 randomly sampled model states $\mathbf{X}=[\mathbf{B};\mathbf{A}]$.
Table E.1: Statistics of the expected norm of stochastic gradients across clients.
Client ID Number of samples Rank Min Max
1 4580 4 0.1286 0.9499
2 4216 19 0.1284 0.7069
3 4873 9 0.1237 0.5774
4 5124 32 0.1499 0.8066
As we can see from Table E.1, the expected norm, i.e., $\mathbb{E}_{\xi\sim\mathcal{D}_{i},\mathbf{S}\sim\mathcal{S}_{i}}\|\nabla_{\mathbf{X}}\tilde{\ell}(\mathbf{X},\xi;\mathbf{S})\|$, is in a moderate range. Notably, the variance is upper-bounded by the expected squared gradient norm:
$\mathbb{E}_{\xi\sim\mathcal{D}_{i},\mathbf{S}\sim\mathcal{S}_{i}}\|\nabla_{\mathbf{X}}\tilde{\ell}(\mathbf{X},\xi;\mathbf{S})-\nabla_{\mathbf{X}}f_{i}^{\mathcal{S}}(\mathbf{X})\|^{2}\leq\mathbb{E}_{\xi\sim\mathcal{D}_{i},\mathbf{S}\sim\mathcal{S}_{i}}\|\nabla_{\mathbf{X}}\tilde{\ell}(\mathbf{X},\xi;\mathbf{S})\|^{2}.$
Therefore, it is generally not hard to find a $\sigma$ and $\rho$ to make Assumption 4.2 hold.
On the other hand, we have
$\left\|\nabla_{\bf X}f_{i}^{\mathcal{S}}({\bf X})\!-\!\nabla_{\bf X}f^{\mathcal{S}}(\bf X)\right\|^{2}$ (7)
$\leq$ $2\left\|\nabla_{\bf X}f_{i}^{\mathcal{S}}({\bf X})\right\|^{2}+2\left\|\nabla_{\bf X}f^{\mathcal{S}}(\bf X)\right\|^{2}$
$\leq$ $2\left\|\nabla_{\bf X}f_{i}^{\mathcal{S}}({\bf X})\right\|^{2}+2\frac{1}{N}\sum_{i=1}^{N}\left\|\nabla_{\bf X}f_{i}^{\mathcal{S}}({\bf X})\right\|^{2}$
$\leq$ $2\mathbb{E}_{\xi\sim\mathcal{D}_{i},\mathbf{S}\sim\mathcal{S}_{i}}\|\nabla_{\mathbf{X}}\tilde{\ell}(\mathbf{X},\xi;\mathbf{S})\|^{2}+2\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{\xi\sim\mathcal{D}_{i},\mathbf{S}\sim\mathcal{S}_{i}}\|\nabla_{\mathbf{X}}\tilde{\ell}(\mathbf{X},\xi;\mathbf{S})\|^{2},$
where the first inequality follows Cauchy-Schwarz inequality, while the last inequality follows Jensen’s inequality. Thus, the deviation can be controlled by the expected gradient norm, which we empirically found to be moderate (see Table E.1). Hence, it is reasonable to impose an upper bound on $\left\|\nabla_{\bf X}f_{i}^{\mathcal{S}}({\bf X})\!-\!\nabla_{\bf X}f^{\mathcal{S}}(\bf X)\right\|^{2}$ as in Assumption 4.3.
@@ Appendix E Empirical Validation of Assumptions
Appendix F Evaluation under Broader Heterogeneity and Increased Number of Clients
To accommodate a larger number of clients, we extend FSLoRA (Algorithm 1) to support partial client participation. Specifically, at each round, the server samples a subset of clients, distributes the current global LoRA modules to them, and aggregates only the updates from these clients. Throughout this section, we fix the partial participation size to 10, i.e., 10 clients are sampled in each round.
F.1 Increasing Resource Heterogeneity and the Number of Clients
@@
We extend our experiments on LLaMA-3.2-3B with the commonsense reasoning benchmark to $50$ clients. We adopt Dirichlet-based partitioning for dataset splits. Specifically, the commonsense reasoning benchmark includes $8$ tasks, and we partitioned them based on the Dirichlet distribution to construct task heterogeneity among $50$ clients. The Dirichlet concentration parameter is set to $\alpha=0.1$. We simulate client resource heterogeneity via different LoRA rank distributions (beyond the limited sketching ratio considered in Section 5). More capable clients are assigned higher ranks, reflecting varying compute capacities. We consider two different rank distributions: normal and heavy-tail distributions in the range $[4,64]$.
@@
Normal distribution: Ranks are sampled from a normal distribution with mean $\mu=\frac{a+b}{2}$ and standard deviation $\sigma=\frac{b-a}{6}$, where $a=4$ and $b=64$. This models a balanced distribution of client capabilities centered around the middle of the range.
@@
Heavy-tail distribution: We sample ranks using an inverse log-normal distribution. Specifically, we draw $x_{i}\sim\text{LogNormal}(\mu,\sigma)$ with $\mu=\log\left(\frac{a+b}{4}\right)$ and $\sigma=1.0$, then set $k_{i}=1/x_{i}$ and apply min-max normalization to scale values into the range $[a,b]$. This results in a heavy-tailed distribution where most clients receive low ranks, reflecting a scenario with many low-capability clients and a few high-capability ones.
@@
Table F.1: Accuracy comparison under different client heterogeneity settings. FSLoRA outperforms baseline methods across both normal and heavy-tail LoRA rank distributions.
@@
As shown in Table F.1, FSLoRA outperforms other methods under both heterogeneity settings. As we move from normal to heavy-tail, where more clients are low-resource, overall performance decreases for all methods. However, FSLoRA exhibits the smallest drop, demonstrating stronger robustness to extreme client heterogeneity.
In Figure 5, we compare the convergence behavior of FSLoRA and three baseline methods under the aforementioned two types of client heterogeneity. Under the normal distribution, FlexLoRA exhibits fast initial progress but falls behind FSLoRA in final accuracy, likely due to approximation errors introduced by truncated SVD. This issue is exacerbated in the heavy-tail distribution, where low-rank clients dominate and SVD truncation causes greater distortion, further degrading FlexLoRA’s performance. Similarly, HeteroLoRA’s reliance on zero-padding reduces optimization efficiency, preventing it from achieving higher accuracy. FLoRA fails to show steady improvement as communication progresses. One potential reason is that frequent model merging and random reinitialization of LoRA modules in each round disrupt the convergence continuity. In contrast, FSLoRA demonstrates robust and stable convergence across both scenarios, achieving the highest overall accuracy.
@@
F.2 Further Increasing the Number of Clients
We further evaluated the performance of FSLoRA by increasing the number of clients to $100$. The results are presented in Table F.2. In this setting, local ranks follow a heavy-tailed distribution as described in the previous subsection, and all other experimental configurations remain unchanged. As shown in the table, FSLoRA maintains its advantage in terms of the average performance when scaling to more clients.
Table F.2: Accuracy comparison when the number of clients is $N=100$. FSLoRA maintains its advantage in terms of the average accuracy.